OpenStudy (anonymous):

Solve the system by graphing. Help Please? y=3x+6 y=-3x

OpenStudy (anonymous):

it is easier to solve by writing \[-3x=3x+6\] and solving for \(x\)

OpenStudy (anonymous):

you get \(-6x=6\) and so \(x=-1\) then if \(x=-1\) you have \(y=-3\times -1=3\)

OpenStudy (compassionate):

I disagree, Satellite, it's easier to solve by graphing it in basic y = mx + b method

OpenStudy (anonymous):

i already have the first one, i just don't know how to do the second one. And I'd wrather just stick with y=mx+b Sorry(:

OpenStudy (anonymous):

@Compassionate do you know how to solve the second one? I don't know what to do because there is only m

OpenStudy (compassionate):

There is no y-intercept for the second on. So your solution will lie cross the y-axis only. That is, go up 3, and over 1, then graph.

OpenStudy (anonymous):

not sure what you mean by :"second one" the answer is not a line, it is a point, namely \((-1,3)\)

OpenStudy (compassionate):

Only given the slope = only graph the slope. It should be a straight line.

OpenStudy (compassionate):

Right, but there is no y-intercept, Satellite. So you're just graphing (-3, 1)

OpenStudy (compassionate):

\[- \frac{ 3 }{ x } = - \frac{ 3 }{ 1 } = \frac{ -3 }{ 1 }\]

OpenStudy (anonymous):

so, i did what you said but i got parallel lines. @satellite73

OpenStudy (anonymous):

y=mx+c is straight line graph equation. 3 is gradient so it steep by 3. 6 is y intercept

OpenStudy (anonymous):

yes, but then the second equation is also a straight line, and there is no solution because you get parallel lines right?

OpenStudy (anonymous):



OpenStudy (anonymous):

yes, exactly what i got. but for the second equation it should be a parallel line to the one you just drew, therefore no solution.

OpenStudy (anonymous):

right? @jimswig88

OpenStudy (anonymous):

y=-3x is same gradient sloping opposite to that line cutting through origin 0,0

OpenStudy (anonymous):

i have (1,-3) for the second line.

OpenStudy (anonymous):

@jimswig88

OpenStudy (anonymous):



OpenStudy (anonymous):

okay. Thanks!!

OpenStudy (anonymous):

yes both lines intercept at -1,3